# MIT-CTP/5776

### v1

@@ Abstract
We introduce the complex Liouville string, a solvable string theory defined by coupling two Liouville theories with complex conjugate central charges $c\in 13+i\mathbb{R}$ on the worldsheet. We compute its amplitudes from first principles and establish a duality with a double-scaled two-matrix integral. We also analyze general worldsheet boundaries and non-perturbative effects in the genus expansion. This theory is capable of capturing aspects of de Sitter quantum gravity in both two and three dimensions, which leads to a precise version of de Sitter holography.
@@ I Introduction
We refer to this theory as the complex Liouville string ($\mathbb{C}$LS). This theory is very interesting from a variety of perspectives and holds a surprising number of lessons in stock. The string amplitudes can be computed exactly from first principles by employing analytic bootstrap techniques. It admits a dual description in terms of a random two-matrix model, and moreover describes a theory of 2d quantum gravity with vacua of both positive and negative cosmological constants. Thus it provides an exact version of holography involving de Sitter (dS) spacetimes, which is reflected in several unusual properties of the worldsheet theory and of the matrix model. The theory allows us to consider observables with general boundaries from the worldsheet, and to study (doubly) non-perturbative effects in the sum over topologies in detail. Finally, the string amplitudes of (1) can also be reinterpreted as cosmological correlators of 3d pure de Sitter gravity with non-dynamical massive particles. This interpretation gives a precise version of a de Sitter/ matrix model correspondence. In this letter, we give an overview of these aspects of $\mathbb{C}$LS and refer to the papers [7, 8, 9, 10] for further details.
@@ III Matrix model
$\omega_{g,n}(\boldsymbol{z})=\int_{0}^{\infty}\prod_{j=1}^{n}\frac{2\pi ip_{j}\text{d}p_{j}\exp(2\pi ip_{j}\sqrt{z_{j}})\text{d}z_{j}}{\sqrt{z_{j}}}\,\mathsf{A}_{g,n}^{(b)}(\boldsymbol{p})~{},$ (12)
@@ IV Worldsheet boundaries and non-perturbative effects
V De Sitter quantum gravity
@@
$S_{\mathrm{dil}}[\Phi,g]=\frac{i}{2b^{2}}\int_{\Sigma_{g,n}}\!\!\!\text{d}^{2}x\sqrt{g}\,\big{(}\Phi\mathcal{R}-\pi^{-1}\sin(2\pi\Phi)\big{)}~{},$ (17)
@@
somewhat reminiscent of [45]. Here $\xi_{j}$ are the locations of the vertex operator insertions. The dS solutions necessitate a (,)(-,-)( - , - ) signature of the metric, similar to the dS saddles of [32]. $m$ can be identified with the color $m_{v}$ appearing in the exact answer (13), which suggests that only the saddles $m\geq 1$ contribute to the path integral (18). (13) also suggests that we should consider more general saddles consisting of degenerated Riemann surfaces connected at nodes whose different components $v\in\mathcal{V}$ can be in different vacua $m_{v}\in\mathbb{Z}_{>0}$. These represent transition amplitudes between universes with different cosmological constants. Evaluating the path integral (18) beyond the leading saddle we uncover the same fluctuation theory as in the Virasoro minimal string [5]. Together this provides a path integral interpretation of the sum over stable graphs (13).
Liouville as the wavefunction of the universe.
The string amplitudes can be reinterpreted as cosmological correlators in $\mathrm{dS}_{3}$ quantum gravity. The Hilbert space of $\mathrm{dS}_{3}$ quantum gravity can be obtained via its relation with $\text{SL}(2,\mathbb{C})$ Chern-Simons theory. Canonical quantization of the phase space on the initial value surface $\Sigma_{g,n}$ with a suitable gravitational polarization as in [46] yields a Hilbert space $\mathcal{H}_{g,n}^{(b)}(\boldsymbol{p})$ spanned by products of left- and right-moving conformal blocks of central charge $c\in 13+i\mathbb{R}_{+}$ and conformal weights $h_{j}\in\frac{1}{2}+i\mathbb{R}$ corresponding to principal series representations of $\text{SL}(2,\mathbb{C})$. The imaginary part of the central charge is set by the dS radius in Planck units [47], $\mathop{\text{Im}}c=\frac{3\ell_{\text{dS}}}{2G_{\text{N}}}$. Conformal weights are related to the masses of massive particles via the standard holographic dictionary, $h=\frac{1}{2}\pm\frac{1}{2}\sqrt{1-m^{2}\ell_{\text{dS}}^{2}}$ [48]. In gravity, we also gauge large diffeomorphisms, thus wavefunctions $\Psi\in\mathcal{H}_{g,n}^{(b)}(\boldsymbol{p})$ must be crossing symmetric in the sense of 2d CFT. Thus wavefunctions transform precisely as 2d CFT correlators with central charge $c\in 13+i\mathbb{R}$ under diffeomorphisms and Weyl transformations.
For de Sitter quantum gravity, we have to specify the wavefunction of the universe $\Psi\in\mathcal{H}_{g,n}^{(b)}(\boldsymbol{p})$ living at future infinity $\mathcal{I}^{+}$. There is no clear principle telling us what $\Psi$ should be. In particular, the analogue of the Hartle-Hawking state [49] neither seems to be well-defined since the sum over topologies diverges badly, nor does it seem to be normalizable [50, 51]. Instead, one can motivate the Liouville correlator $\Psi=\langle V_{p_{1}}\cdots V_{p_{n}}\rangle$ to be a valid wavefunction of the universe. It can be uniquely bootstrapped by requiring normalizability of the wavefunction and appropriate factorization under degeneration of $\mathcal{I}^{+}$.
String amplitudes as cosmological correlators.
Given $\Psi_{g,n}\in\mathcal{H}_{g,n}^{(b)}$, we can compute cosmological correlators of non-dynamical massive particles by integrating over the metrics of $\mathcal{I}^{+}$ [52],
$\lVert\Psi_{g,n}\rVert^{2}=\int_{\mathcal{M}_{g,n}}|\Psi_{g,n}|^{2}~{}.$ (20)
This recovers the amplitudes (2) up to normalization.
de Sitter/matrix model correspondence.
$\mathsf{A}_{g,n}^{(b)}$ computes cosmological correlators of massive particles in a specific state of the universe. Given the correspondence with the matrix model, this constitutes a precise version of holography where the bulk observables are cosmological correlators. Since we already integrate over all metrics in (20) and topology fluctuates in quantum gravity, we should actually consider the genus resummed amplitudes (3) as the true observables of the bulk theory. The parameter $S_{0}$ is a stand in for $\log N$, where $N$ measures the degrees of freedom in the matrix model, and thus plays the role of the de Sitter entropy.
VI Acknowledgements
@@
We would like to thank dio Anninos, Aleksandr Artemev, Mattia Biancotto, Matthias R. Gaberdiel, Davide Gaiotto, Alessandro Giacchetto, Victor Godet, Victor Gorbenko, Clifford V. Johnson, Shota Komatsu, Adam Levine, Raghu Mahajan, Juan Maldacena, Alex Maloney, Marcos Mariño, Nico Valdes-Meller, Mehrdad Mirbabayi, Miguel Montero, Nikita Nekrasov, Hirosi Ooguri, Boris Post, Maximilian Schwick, Ashoke Sen, Douglas Stanford, Jörg Teschner, Cumrun Vafa, Herman Verlinde, Edward Witten and Xi Yin for discussions. SC, LE and BM thank l’Institut Pascal at Université Paris-Saclay, with the support of the program “Investissements d’avenir” ANR-11-IDEX-0003-01, and SC and VAR thank the Kavli Institute for Theoretical Physics (KITP), which is supported in part by grant NSF PHY-2309135, for hospitality during the course of this work. VAR is supported in part by the Simons Foundation Grant No. 488653, by the Future Faculty in the Physical Sciences Fellowship at Princeton University, and a DeBenedictis Postdoctoral Fellowship and funds from UCSB. SC is supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567 (High Energy Theory research), DOE Early Career Award DE-SC0021886 and the Packard Foundation Award in Quantum Black Holes and Quantum Computation. BM gratefully acknowledges funding provided by the Sivian Fund at the Institute for Advanced Study and the National Science Foundation.

### v2

@@ Abstract
We introduce the complex Liouville string, a solvable string theory defined by coupling two Liouville theories with complex conjugate central charges $c\in 13+i\mathbb{R}$ on the worldsheet. We compute its amplitudes from first principles and establish a duality with a double-scaled two-matrix integral. We also analyze general worldsheet boundaries and non-perturbative effects in the genus expansion. By expressing the complex Liouville string as a 2d dilaton gravity theory with a sine potential, we show that it admits both AdS2 and dS2 vacua.
@@ I Introduction
We refer to this theory as the complex Liouville string ($\mathbb{C}$LS). This theory is very interesting from a variety of perspectives and holds a surprising number of lessons in stock. The string amplitudes can be computed exactly from first principles by employing analytic bootstrap techniques. It admits a dual description in terms of a random two-matrix model, and moreover describes a theory of 2d quantum gravity with vacua of both positive and negative cosmological constants. Thus it provides an exact version of holography involving de Sitter (dS) spacetimes, which is reflected in several unusual properties of the worldsheet theory and of the matrix model. The theory allows us to consider observables with general boundaries from the worldsheet, and to study (doubly) non-perturbative effects in the sum over topologies in detail. In this letter, we give an overview of these aspects of $\mathbb{C}$LS and refer to the papers [7, 8, 9, 10] for further details.
@@ III Matrix model
$\omega_{g,n}(\boldsymbol{z})=\int_{0}^{\infty}\prod_{j=1}^{n}\frac{-2\pi ip_{j}\text{d}p_{j}\exp(2\pi ip_{j}\sqrt{z_{j}})\text{d}z_{j}}{\sqrt{z_{j}}}\,\mathsf{A}_{g,n}^{(b)}(\boldsymbol{p})~{},$ (12)
@@ IV Worldsheet boundaries and non-perturbative effects
V Gravitational path integral
@@
$S_{\mathrm{dil}}[\Phi,g]=\frac{i}{2b^{2}}\int_{\Sigma_{g,n}}\!\!\!\text{d}^{2}x\sqrt{g}\,\big{(}\Phi\mathcal{R}+\frac{1}{\pi}\sin(2\pi\Phi)\big{)}~{},$ (17)
@@
somewhat reminiscent of [45, 46]. Here $\xi_{j}$ are the locations of the vertex operator insertions. We have thus found an infinite number of classical saddles with constant dilaton labelled by an integer $m\in\mathbb{Z}$. For even $m$, the Ricci scalar is negative, whereas for odd $m$ we obtain a positive scalar curvature. The sine dilaton gravity theory (17) thus admits a family of alternating anti-de Sitter and de Sitter vacua. The dS solutions necessitate a (,)(-,-)( - , - ) signature of the metric, similar to the dS saddles of [32]. $m$ can be identified with the color $m_{v}$ appearing in the exact answer (13), which suggests that only the saddles $m\geq 1$ contribute to the path integral (18). (13) also suggests that we should consider more general saddles consisting of degenerated Riemann surfaces connected at nodes whose different components $v\in\mathcal{V}$ can be in different vacua $m_{v}\in\mathbb{Z}_{>0}$. These represent transition amplitudes between universes with different cosmological constants. Evaluating the path integral (18) beyond the leading saddle we uncover the same fluctuation theory as in the Virasoro minimal string [5]. Together this provides a path integral interpretation of the sum over stable graphs (13).
VI Outlook
We conclude with an outlook on the relation of the complex Liouville string and 3d de Sitter quantum gravity.
In [47] we carefully quantize the gravitational phase space of dS3 quantum gravity on a compact, hyperbolic Cauchy slice $\Sigma_{g,n}$. We motivate that the Liouville correlator $\Psi^{(b)}_{g,n}=\langle V_{p_{1}}\cdots V_{p_{n}}\rangle_{\Sigma_{g}}$ with central charge $c\in 13+i\mathbb{R}$ is a natural wavefunction of the universe at $\mathcal{I}^{+}$. It corresponds to an expanding spacetime with a big bang singularity in the past, and its norm is precisely computed by the string amplitude $\mathsf{A}^{(b)}_{g,n}$ of the complex Liouville string. We interpret the norm of the wavefunction as defining cosmological correlators of massive particles in dS3, integrated over the metric at $\mathcal{I}^{+}$. This constitutes a first piece of evidence toward a microscopic realization of dS3 in terms of the matrix integral dual of the complex Liouville string. The second piece of evidence ties to the Gibbons-Hawking de Sitter entropy conjecture [48, 49]. According to this, the cosmological event horizon surrounding an observer in dS captures an entropy, which is computed by the logarithm of the Euclidean gravity path integral on the sphere (the Euclidean continuation of de Sitter). We observe in [47] that truncating the eigenvalue density (11) at its first zero, the Gibbons-Hawking entropy can be reproduced microscopically by counting the number of eigenvalues in the dual matrix model.
VII Acknowledgements
@@
We would like to thank Dionysios Anninos, Aleksandr Artemev, Mattia Biancotto, Matthias R. Gaberdiel, Davide Gaiotto, Alessandro Giacchetto, Victor Godet, Victor Gorbenko, Clifford V. Johnson, Shota Komatsu, Adam Levine, Raghu Mahajan, Juan Maldacena, Alex Maloney, Marcos Mariño, Nico Valdes-Meller, Mehrdad Mirbabayi, Miguel Montero, Nikita Nekrasov, Hirosi Ooguri, Boris Post, Maximilian Schwick, Ashoke Sen, Douglas Stanford, Jörg Teschner, Cumrun Vafa, Herman Verlinde, Edward Witten and Xi Yin for discussions. SC, LE and BM thank l’Institut Pascal at Université Paris-Saclay, with the support of the program “Investissements d’avenir” ANR-11-IDEX-0003-01, and SC and VAR thank the Kavli Institute for Theoretical Physics (KITP), which is supported in part by grant NSF PHY-2309135, for hospitality during the course of this work. VAR is supported in part by the Simons Foundation Grant No. 488653, by the Future Faculty in the Physical Sciences Fellowship at Princeton University, and a DeBenedictis Postdoctoral Fellowship and funds from UCSB. SC is supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567 (High Energy Theory research), DOE Early Career Award DE-SC0021886 and the Packard Foundation Award in Quantum Black Holes and Quantum Computation. BM gratefully acknowledges funding provided by the Sivian Fund at the Institute for Advanced Study and the National Science Foundation with grant number PHY-2207584.